Please solve this equation for x :
1/X - 4X = 7.5

Answer by stanbon(75887)

1/X - 4X = 7.5
1/x - 4x = 15/2
Multiply thru by 2x to get:
2 - 8x^2 = 15x
8x^2 + 15x -2 = 0
x = [-15 +- sqrt(225-4*8*-2)]/16
x = [-15 +- sqrt(289)]/16
x = [-15 +- 17]/16
x = -32/16 or 2/16
x = -2 or x = 1/8
=======================
Cheers,
Stan H.

=================================